Sodium trioxalatocobaltate (III) trihydrate is prepared by thefollowing reactions:
   [Co(H2O)6]Cl2 +K2C2O4 x H2O--> CoC2O4 + 2KCl +7H2O
2CoC2O4 + 4H2O +H2O2 +4Na2C2O4 -->2Na3[Co(C2O4)3] x3H2O + 2NaOH

What is the percent yield ofNa3[Co(C2O4)3] x3H2O if 7.6 g is obtained from 12.5 g of[Co(H2O)6]Cl2?
Answer #1
[Co(H2O)6]Cl2 +K2C2O4 x H2O--> CoC2O4 + 2KCl +7H2O
2CoC2O4 + 4H2O +H2O2 +4Na2C2O4 -->2Na3[Co(C2O4)3] x3H2O + 2NaOH

molar mass of [Co(H2O)6]Cl2 = 59 +6(18) + 2(35.5) = 238 g/mol
n([Co(H2O)6]Cl2) = 12.5 g /238gmol-1
                            = 0.0525 mol
n(CoC2O4) =n([Co(H2O)6]Cl2 * 1mol ofCoC2O4 / 1mol ofn([Co(H2O)6]Cl2
                  = 0.0525 mol
n(Na3[Co(C2O4)3]) =n(CoC2O4) * 2mol ofNa3[Co(C2O4)3] / 2molof CoC2O4
                                         
= 0.0525 mol

Molar mass ofNa3[Co(C2O4)3] = 3(23)+ 59 + 3(2*12+4*16)
                                                 = 392 g/mol
Theorically mass ofNa3[Co(C2O4)3] produced= 392 g/mol * 0.0525 mol
                                                                        = 20.502g
Percentage yield = 7.6 / 20.502 * 100%
                        = 37.069%